OpenStudy (anonymous):

Medal! Trevor solved the system of equations below. What mistake did he make in his work? 2x + y = 5 x − 2y = 10 y = 5 − 2x x − 2(5 − 2x) = 10 x − 10 + 4x = 10 5x − 10 = 10 5x = 0 x = 0 2(0) + y = 5 y = 5 He should have substituted 5 + 2x He combined like terms incorrectly, it should have been 4x instead of 5x He subtracted 10 from the right side instead of adding 10 to the right side He made no mistake

OpenStudy (anonymous):

Which do you think is the correct answer? Any ideas?

OpenStudy (dayakar):

5x-10=10 5x= 10+10 5x= 20

OpenStudy (anonymous):

i would like to learn it and then find the answer.

OpenStudy (dayakar):

x= 20/4 x=5

OpenStudy (dayakar):

sorry x=20/5 x=4

OpenStudy (anonymous):

im very confuse

OpenStudy (anonymous):

Are there any answers from the multiple choice ones that you can eliminate? For example, look at A: He should have substituted 5+2x for y Should he have done this, or would this have been an incorrect step?

OpenStudy (dayakar):

u did the mistake in fifth line

OpenStudy (anonymous):

A) He should have substituted 5 + 2x B)He combined like terms incorrectly, it should have been 4x instead of 5x C)He subtracted 10 from the right side instead of adding 10 to the right side D) He made no mistake

OpenStudy (dayakar):

5x-10 =10 next line is wrong

OpenStudy (anonymous):

?

OpenStudy (dayakar):

add both sides 10 what do u get

OpenStudy (anonymous):

20

OpenStudy (dayakar):

5x-10+10=10+10 5x= 20 is it clear

OpenStudy (mathmale):

@Buttercup214 : I'd suggest, seriously, that you working on solving this system of equations yourself. If you can do that correctly (and this is a skill you must learn), you should have no trouble finding Trevor's mistake. 2x + y = 5 x − 2y = 10

OpenStudy (anonymous):

um ok

OpenStudy (mathmale):

Have you begun your solution yet? You can always ask clarifying questions if you like.

OpenStudy (anonymous):

its ok i found the answer.